OpenStudy (anonymous):

3x^2+16x-44=0 solve by factoring

OpenStudy (jack1):

(3x + something) (x - something)

OpenStudy (jack1):

the somethings you work out by (what x what = -44 and what + what = +16)

OpenStudy (jack1):

remember there's a 3x in there though

OpenStudy (jack1):

factors of 44 are 1,44; 2,22 ; 4,11

OpenStudy (jack1):

so pick which pair fits your equation

OpenStudy (anonymous):

@ashley007

OpenStudy (anonymous):

2 22?

OpenStudy (jack1):

yep, as 3x 2 is 6, and 22-6 = 16 so the 2 will be the -something and the 22 the + something

OpenStudy (anonymous):

thats final answe?

OpenStudy (jack1):

no, that's how you factor it

OpenStudy (jack1):

so plug that into you equation and you get (3x +22) ( x-2) = 0 therefore either the 3x +22 = 0 or the x-2 = 0

OpenStudy (jack1):

so rearrange those 2 equations to see what x = in each, and those are your answers for the x-intercepts

OpenStudy (anonymous):

Sorry find 2 factors of 44*3 which add upto 16...

OpenStudy (anonymous):

hmm 7 and 9

OpenStudy (anonymous):

@higherlearning ....

OpenStudy (anonymous):

i dont get that

OpenStudy (anonymous):

Fyn multiply 44 by 3.. Wat do u get?

OpenStudy (anonymous):

132

OpenStudy (anonymous):

Good. Now find any 2 factors of 132 which mudt add upto 16....

OpenStudy (anonymous):

@higherlearning ?!??

OpenStudy (anonymous):

8

OpenStudy (anonymous):

I'll help u out... Its 22 &6...now can u do it???

OpenStudy (anonymous):

how do those add up to 16

OpenStudy (anonymous):

Sorry dubtract it... I'm really tired....

OpenStudy (anonymous):

I bet u can do it now... I dtart... U continue.... 3x^2+16x-44=0 3x^2+22x-6x-44=0 3x(x-2)+22(x-2)=0 now continue @higherlearning ...

OpenStudy (anonymous):

@higherlearning ??? R u there ??

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Comeon u can do it now... Try.. I'll correct u..

OpenStudy (anonymous):

Now we get 2 factors.. (3x+22) &(x-2)... Solve them 1 by 1... 3x+22=0 3x=-22 X=-22/3.. similarly do d second one...

OpenStudy (anonymous):

(x-2)=0 x=2 Therefore x=-22/3,2 !!! There u go @higherlearning !!